a store stocks 50 birthday cards, 45 get well cards and 10 anniversary cards

what simplest form in a fraction is the get well cards
1
Expert's answer
2011-11-22T09:28:20-0500
total number of cards
50+45+10=105
hence fractional part of get well cards is& 45/105=9/21=3/7&